Relative cost random access machines

Martti Penttonen

Report A/1998/5

ISBN 951-781-607-3 ISSN 0787-6416

UNIVERSITY OF KUOPIO Department of Computer Science

and Applied Mathematics

P.O.Box 1627, FIN-70211 Kuopio, FINLAND